# Modeling and Analysis of Land-to-Ship Maritime Wireless Channels at 5.8 GHz

### v1

@@ IV SWIFT Fading
where $T_{i}$ and $\lambda_{i}$ are the fluctuation period and wavelength of the $i$-th composite wave, respectively, $\xi_{i}$ denotes the initial phase of the $i$-th composite wave that follows a uniform random distribution in $[0,2\pi)$. $A_{i}$ represents the amplitude of the $i$-th composite wave, which depends on the frequency spectrum of sea waves $S(\omega)$, and the classic Pierson-Moskowitz (P-M) spectrum is often utilized to model $S(\omega)$ for low computational complexity [15]
$A_{i}=\sqrt{2S(\omega_{i})\Delta\omega},\quad S_{\text{P-M}}(\omega)=\frac{a_{0}g^{2}}{\omega^{5}}e^{-\beta\left(\frac{g}{v_{\text{w}}\omega}\right)^{4}}$ (16)
[…]
@@
For the rotational motion of the Rx antenna, also as shown in Fig. 6, we define three types of rotational motion, i.e., roll, pitch, and yaw, around the x, y, and z coordinate axes, respectively. With the Rx antenna center as the origin and rotation center, and the vessel’s forward direction as the x-axis direction, these three rotational motions can be described by the rotation matrices in (20), where we assume that the roll angel $\phi$, pitch angle $\theta$, and yaw angle $\psi$ follow sinusoidal motions as $\phi(t)=\Phi\sin(\omega_{\phi}t+\alpha_{\phi}),\theta(t)=\Theta\sin(\omega_{\theta}t+\alpha_{\theta}),\psi(t)=\Psi\sin(\omega_{\psi}t+\alpha_{\psi})$ where $\Phi$, $\Theta$, and $\Psi$ are the maximum rotation angles, $\omega_{\phi}$, $\omega_{\theta}$, and $\omega_{\psi}$ are the frequencies of the rotation angle variations, $\alpha_{\phi}$, $\alpha_{\theta}$, and $\alpha_{\psi}$ follow a uniform random distribution in $[0,2\pi]$.
Considering that the shore-to-ship direction aligns with the x-axis, the unit LoS vector $\mathbf{u}_{0}$ is $[\cos\alpha_{0},0,\sin\alpha_{0}]$, where $\alpha_{0}=\arctan\frac{h_{t}-h_{r}}{d}$. The direction vector of the Rx antenna after rotation is $\mathbf{u}=\mathbf{R}_{z}(\psi)\mathbf{R}_{y}(\theta)\mathbf{R}_{x}(\phi)\mathbf{u}_{0}=[u_{x},u_{y},u_{z}]$. For an omnidirectional antenna, we only consider the impact of elevation angle changes on antenna gain. The Rx antenna elevation angle relative to the horizontal plane is $\arcsin(u_{z})$, while the antenna gain loss due to antenna pattern rotation is $L_{\text{g}}=20\log_{10}(F(\arcsin(u_{z})))$, where $F(\cdot)$ is the radiation pattern of the omnidirectional antenna, and $u_{z}=-\sin(\theta)\cos(\alpha_{0})+\cos(\theta)\cos(\phi)\sin(\alpha_{0})$. For polarization mismatch loss $L_{p}$, since both Tx and Rx antennas are vertically polarized, we consider the inner product of polarization vectors as the polarization mismatch loss. Defining the Tx and Rx polarization vectors $\mathbf{p}_{t}$ and $\mathbf{p}_{r}$ as $[0,0,1]^{T}$, the inner product of polarization vectors after rotation is $\rho=|\mathbf{p}_{t}^{T}\mathbf{R}_{z}(\psi)\mathbf{R}_{y}(\theta)\mathbf{R}_{x}(\phi)\mathbf{p}_{r}|=|\cos(\theta)\cos(\phi)|$, and $L_{\text{p}}=20\log_{10}(\rho)$.
@@
Figure 7: Measured and simulated results of SWIFT fading.
Remark 1: As can be observed from Fig. 7, signal variation is more severe for high Rx antennas, attributed to the fact that the amplitude of swaying and fluctuation is more acute for a higher Rx antenna. The proposed SWIFT fading model can well capture the fading intensity changing trend as the antenna height and sea states vary. On the other hand, the simulated signal strength is generally more concentrated than the measured ones, since (15) and the vessel motion models may be simplified compared with the actual circumstances.
@@ V Small-Scale Fading
In marine environments within $d_{\text{LoS}}^{\text{vision}}$, propagation is dominated by direct paths and sea-surface reflections due to sparse obstacles. Additionally, the rough sea surface creates multiple scattered paths, effectively acting as fragmented mirrors. These conditions make the TWDP model suitable within LoS distances, particularly under adverse sea conditions [3]. The mathematical expression of the TWDP model is [41]
$\tilde{V}=V_{1}e^{j\psi_{1}}+V_{2}e^{j\psi_{2}}+\tilde{V}_{\text{dif}},$ (21)
where $\tilde{V}$ is the received baseband signal, $V_{1}$ and $V_{2}$ (0 $<$ $V_{2}$ $<$ $V_{1}$) represent the constant amplitudes of two dominant specular multipath components, with their phases $\psi_{1}$ and $\psi_{2}$ independently and uniformly distributed over the range $[0,2\pi)$. The diffuse component $\tilde{V}_{\text{dif}}$ is modeled as a complex Gaussian random variable with zero mean and variance $2\sigma^{2}$, representing the combined contribution of numerous weaker multipath waves. In the case where $V_{2}=0$, the TWDP model simplifies to the Rician distribution. When both $V_{1}$ and $V_{2}$ are zero, the TWDP model further degenerates to the Rayleigh distribution. The PDF of the TWDP model is given by [41]
$f_{U}(u)$ $=\frac{u}{\pi\sigma^{2}}\exp\left\{-\frac{u^{2}}{2\sigma^{2}}-B_{1}\right\}$ (22)
×$\times\int_{0}^{\pi}\exp\{B_{2}B_{1}\cos x\}I_{0}\left(\frac{u}{\sigma}\sqrt{2B_{1}(1-\Delta\cos x)}\right)\,dx,$
@@ VI Channel Sparsity
In maritime communications, the sparse distribution of scatterers at sea incurs significant sparsity in the wireless channel[3]. While there exist a few potential metrics to evaluate channel sparsity, such as the number of MPCs, channel degrees of freedom [42, 43], Gini index[44], and Rician $K$ factor [45], the Gini index and Rician $K$ factor have been shown to be the most reasonable choices [46]. We thus employ these two measures herein. The Rician $K$ factor and Gini index $G$ are respectively defined as [45, 46]
@@
where $P_{\text{tot}}$ is the total power of all MPCs in a PDP, $P_{\text{LoS}}$ represents the power of the LoS or the strongest path, $N$ is the total number of MPCs in a PDP, and $P_{n}$ is the power of each MPC arranged in an ascending order, i.e., $P_{1}\leq P_{2}\leq\dots\leq P_{N}$. If the LoS path exists, we usually define that $P_{\text{LoS}}=P_{N}$, where the LoS path possesses the highest power among all MPCs. The Rician $K$ factor represents the ratio of the power of the LoS or the strongest path to the power of the remaining MPCs. A higher $K$ indicates greater channel sparsity. The Gini index measures power concentration across MPCs, which is increasingly being adopted as a standard for measuring channel sparsity [46]. As shown by (26), the value of $G$ ranges from 0 to 1. A larger $G$ value indicates that power is concentrated in fewer MPCs, prompting a higher degree of sparsity. When $G=1$, it suggests that all the power is concentrated in the LoS or strongest path, whereas $G=0$ implies that the power is evenly distributed across all MPCs.666 Note that channels exhibiting high sparsity in terms of the Gini index and Rician $K$ factor imply that few MPCs contain most of the received power, naturally corresponding to a small number of significant MPCs. Unlike the Rician $K$ factor, the Gini index is scale-invariant and depends solely on power proportions, making it a more accurate measure of channel sparsity[44].
@@
Fig. 10 illustrates the Rician $K$ factor versus the Gini index under different fixed-point measurement conditions across all distance ranges, along with least squares fits. The maritime Gini index concentrates in the range of 0.965 to 0.985. As a comparison, the Gini index varies between about 0.4 and 0.8 at 5.9 GHz with a bandwidth of 30 MHz for the urban scenario [45], and between about 0.75 and 0.92 at 6 GHz with a bandwidth of 400 MHz for the indoor office scenario [46]. According to Lemma 1 and Lemma 2, an increase in the bandwidth will raise the Gini index. Therefore, the maritime wireless channel is sparser than typical urban and indoor scenarios. More importantly, over all the fixed-point measurement data, the least squares fitting slope in Fig. 10 suggests that a larger $K$ factor tends to be associated with a higher Gini index, which is consistent with the characteristics of increased channel sparsity.
@@ VIII Conclusions
This paper has presented characterizations of land-to-ship maritime wireless channels at 5.8 GHz. First, we have proposed a new large-scale path loss model, i.e., the dual-slope CI-MTR model, that enjoys solid physical foundation, low complexity, high accuracy, and good adaptability over a vast range of distances and different sea states. It is observed that windy sea environments arouse higher deviation in large-scale path loss. Second, the concept of SWIFT fading has been presented, and the proposed SWIFT fading model can align with measured data, particularly for small movements of the antenna. Third, unlike most terrestrial scenarios where Rician fading can well depict the small-scale fading, asymmetric Laplace distribution is effective under calm sea conditions, while the TWDP model better captures the bimodality of fading behavior under adverse sea conditions. Furthermore, quantitative analysis based on the Gini index implies that the maritime propagation scenario possesses a higher degree of sparsity compared with typical terrestrial counterparts.
The following aspects may be considered for future work:
  • Conducting channel measurements and modeling across diverse sea areas to generalize current findings.
@@
For the minimum resolvable multipath delay interval $\Delta\tau_{2}$, denote with $M$ the number of resolvable MPCs within a time delay bin of width $\Delta\tau_{1}$, then $M\geq 1$. At a certain time instant $t$, the received signal $r_{n}(\tau)$ within the $n$-th delay bin ($n=1,...,N$) of width $\Delta\tau_{1}$ can be expressed as [32]

### v2

@@ IV SWIFT Fading
where $T_{i}$ and $\lambda_{i}$ are the fluctuation period and wavelength of the $i$-th composite wave, respectively, $\xi_{i}$ denotes the initial phase of the $i$-th composite wave that follows a uniform random distribution in $0,2\pi)$. $A_{i}$ represents the amplitude of the $i$-th composite wave, which depends on the frequency spectrum of sea waves $S(\omega)$, and the classic Pierson-Moskowitz (P-M) spectrum is often utilized to model $S(\omega)$ for low computational complexity [15]
$A_{i}=\sqrt{2S(\omega_{i})\Delta\omega},\quad S_{\text{P-M}}(\omega)=\frac{a_{0}g^{2}}{\omega^{5}}\exp[-\beta\left(\frac{g}{v_{\text{w}}\omega}\right)^{4}],$ (16)
where $\omega_{i}$ is the angular frequency with $\omega_{i}=2\pi/T_{i}$, $\Delta\omega$ denotes the frequency interval between adjacent harmonic waves, $a_{0}=8.1\times 10^{-3}$, $\beta=0.74$, $g$ is the gravitational acceleration equal to $9.81~\text{m/s}^{2}$.
@@
For the rotational motion of the Rx antenna, also as shown in Fig. 6, we define three types of rotational motion, i.e., roll, pitch, and yaw, around the x, y, and z coordinate axes, respectively. With the Rx antenna center as the origin and rotation center, and the vessel’s forward direction as the x-axis direction, these three rotational motions can be described by the rotation matrices in (20), where we assume that the roll angel $\phi$, pitch angle $\theta$, and yaw angle $\psi$ follow sinusoidal motions as $\phi(t)=\Phi\sin(\omega_{\phi}t+\alpha_{\phi}),\theta(t)=\Theta\sin(\omega_{\theta}t+\alpha_{\theta}),\psi(t)=\Psi\sin(\omega_{\psi}t+\alpha_{\psi})$ where $\Phi$, $\Theta$, and $\Psi$ are the maximum rotation angles, $\omega_{\phi}$, $\omega_{\theta}$, and $\omega_{\psi}$ are the frequencies of the rotation angle variations, $\alpha_{\phi}$, $\alpha_{\theta}$, and $\alpha_{\psi}$ follow a uniform random distribution in $0,2\pi]$.
Considering that the shore-to-ship direction aligns with the x-axis, the unit LoS vector $\mathbf{u}_{0}$ is $\cos\alpha_{0},0,\sin\alpha_{0}]$, where $\alpha_{0}=\arctan\frac{h_{t}-h_{r}}{d}$. The direction vector of the Rx antenna after rotation is $\mathbf{u}=\mathbf{R}_{z}(\psi)\mathbf{R}_{y}(\theta)\mathbf{R}_{x}(\phi)\mathbf{u}_{0}=[u_{x},u_{y},u_{z}]$. For an omnidirectional antenna, we only consider the impact of elevation angle changes on antenna gain. The Rx antenna elevation angle relative to the horizontal plane is $\arcsin(u_{z})$, while the antenna gain loss due to antenna pattern rotation is $L_{\text{g}}=20\log_{10}(F(\arcsin(u_{z})))$, where $F(\cdot)$ is the radiation pattern of the omnidirectional antenna, and $u_{z}=-\sin(\theta)\cos(\alpha_{0})+\cos(\theta)\cos(\phi)\sin(\alpha_{0})$. For polarization mismatch loss $L_{p}$, since both Tx and Rx antennas are vertically polarized, we consider the inner product of polarization vectors as the polarization mismatch loss. Defining the Tx and Rx polarization vectors $\mathbf{p}_{t}$ and $\mathbf{p}_{r}$ as $0,0,1]^{T}$, the inner product of polarization vectors after rotation is $\rho=|\mathbf{p}_{t}^{T}\mathbf{R}_{z}(\psi)\mathbf{R}_{y}(\theta)\mathbf{R}_{x}(\phi)\mathbf{p}_{r}|=|\cos(\theta)\cos(\phi)|$, and $L_{\text{p}}=20\log_{10}(\rho)$.
@@
Figure 7: Measured data and simulated SWIFT fading based on the MTR model with sea wave fluctuations and antenna motions incorporated.
Remark 2: As can be observed from Fig. 7, both measured and simulated signal variations are more severe for high Rx antennas, attributed to the fact that the amplitude of swaying and fluctuation is more acute for a higher Rx antenna. The proposed SWIFT fading model can well capture the fading intensity changing trend as the antenna height and sea states vary. On the other hand, the simulated signal strength is generally more concentrated than the measured ones, since (15) and the vessel motion models may be simplified compared with the actual circumstances.
It is noteworthy that the performance of the proposed SWIFT fading model might be limited under extreme sea conditions or for larger antenna displacements, as these scenarios could introduce additional propagation dynamics not fully captured by the current formulation.
@@ V Small-Scale Fading
In marine environments within $d_{\text{LoS}}$, propagation is dominated by direct paths and sea-surface reflections due to sparse obstacles. Additionally, the rough sea surface creates multiple scattered paths, effectively acting as fragmented mirrors. These conditions make the TWDP model suitable within LoS distances, particularly under adverse sea conditions [3]. The mathematical expression of the TWDP model is [31]
$\tilde{V}=V_{1}\exp(j\psi_{1})+V_{2}\exp(j\psi_{2})+\tilde{V}_{\text{dif}},$ (21)
where $\tilde{V}$ is the received baseband signal, $V_{1}$ and $V_{2}$ (0 << $V_{2}$ << $V_{1}$) represent the constant amplitudes of two dominant specular multipath components, with their phases $\psi_{1}$ and $\psi_{2}$ independently and uniformly distributed over the range $0,2\pi)$. The diffuse component $\tilde{V}_{\text{dif}}$ is modeled as a complex Gaussian random variable with zero mean and variance $2\sigma^{2}$, representing the combined contribution of numerous weaker multipath waves. In the case where $V_{2}=0$, the TWDP model simplifies to the Rician distribution. When both $V_{1}$ and $V_{2}$ are zero, the TWDP model further degenerates to the Rayleigh distribution. The PDF of the TWDP model is given by [31]
$f_{U}(u)$ $=\frac{u}{\pi\sigma^{2}}\exp\left(-\frac{u^{2}}{2\sigma^{2}}-B_{1}\right)$ (22)
$\cdot\int_{0}^{\pi}\exp\left(B_{1}B_{2}\cos x\right)I_{0}\left[\frac{u\sqrt{2B_{1}(1-\Delta\cos x)}}{\sigma}\right]dx,$
@@ VI Channel Sparsity
In maritime communications, the sparse distribution of scatterers at sea incurs significant sparsity in the wireless channel[3]. While there exist a few potential metrics to evaluate channel sparsity, such as the number of multipath components (MPCs), channel degrees of freedom [43, 44], Gini index[32], and Rician $K$ factor [45], the Gini index and Rician $K$ factor have been shown to be the most reasonable choices [46]. The Rician $K$ factor and Gini index $G$ are respectively defined as [45, 46]
@@
where $P_{\text{tot}}$ is the total power of all MPCs in a PDP, $P_{\text{LoS}}$ represents the power of the LoS or the strongest path, $N$ is the total number of MPCs in a PDP, and $P_{n}$ is the power of each MPC arranged in an ascending order, i.e., $P_{1}\leq P_{2}\leq\dots\leq P_{N}$. If the LoS path exists, we usually define that $P_{\text{LoS}}=P_{N}$, where the LoS path possesses the highest power among all MPCs. The Rician $K$ factor represents the ratio of the power of the LoS or the strongest path to the power of the remaining MPCs. A higher $K$ indicates greater channel sparsity. The Gini index measures power concentration across MPCs, which is increasingly being adopted as a standard for measuring channel sparsity [46]. As shown by (26), the value of $G$ ranges from 0 to 1. A larger $G$ value indicates that power is concentrated in fewer MPCs, prompting a higher degree of sparsity. When $G=1$, it suggests that all the power is concentrated in the LoS or strongest path, whereas $G=0$ implies that the power is evenly distributed across all MPCs. [^6] Note that channels exhibiting high sparsity in terms of the Gini index and Rician $K$ factor imply that few MPCs contain most of the received power, naturally corresponding to a small number of significant MPCs. Unlike the Rician $K$ factor, the Gini index is scale-invariant and depends solely on power proportions, making it a more accurate measure of channel sparsity[32].
@@
Fig. 10 illustrates the Rician $K$ factor versus the Gini index under different fixed-point measurement conditions across all distance ranges, along with least squares fits. Note that we have examined variations of the Rician $K$ factor and the Gini index against distance, but have not observed any obvious trend. Thus, the results presented herein include all measured distances. It is evident from Fig. 10 that the maritime Gini index concentrates in the range of 0.965 to 0.985. As a comparison, the Gini index varies between about 0.4 and 0.8 at 5.9 GHz with a bandwidth of 30 MHz for the urban scenario [45], and between about 0.75 and 0.92 at 6 GHz with a bandwidth of 400 MHz for the indoor office scenario [46]. According to Lemma 1 and Lemma 2, an increase in the bandwidth will raise the Gini index. Therefore, the maritime wireless channel is sparser than typical urban and indoor scenarios. More importantly, over all the fixed-point measurement data, the least squares fitting slope in Fig. 10 suggests that a larger $K$ factor tends to be associated with a higher Gini index, which is consistent with the characteristics of increased channel sparsity. Additionally, the Rician $K$ factor and the Gini index are slightly lower on Day 1 for the same antenna height, indicating that rougher sea conditions are likely to render more diffuse scattering power.
@@ VIII Conclusions
This paper has analyzed land-to-ship maritime wireless channels at 5.8 GHz, uncovering fundamental distinctions from terrestrial channels, such as multi-timescale fading (including SWIFT fading), channel sparsity, and pronounced environmental sensitivity. A series of models have been developed to comprehensively characterize the channel. First, we have proposed a new large-scale path loss model, i.e., the dual-slope CI-MTR model, that enjoys solid physical foundation, low complexity, high accuracy, and good adaptability over a vast range of distances and different sea states. It is observed that windy sea environments arouse higher deviation in large-scale path loss. Second, the concept of SWIFT fading has been presented, and the proposed SWIFT fading model can align with measured data, particularly for small movements of the antenna. Third, unlike most terrestrial scenarios where Rician fading can well depict the small-scale fading, the asymmetric Laplace distribution is effective under calm sea conditions, while the TWDP model better captures the bimodality of fading behavior under adverse sea conditions. Furthermore, quantitative analysis based on the Gini index implies that the maritime propagation scenario possesses a higher degree of sparsity compared with typical terrestrial counterparts. These findings advance the understanding of maritime propagation, offering valuable insights for system design in dynamic oceanic environments.
While the aforementioned models are validated for 5.8 GHz, their generalizability to other frequency bands or maritime environments remains an open question. Additionally, this study is limited to the specific sea states, ship type, and carrier bandwidth. Therefore, the following aspects may be considered for future work:
  • Conducting channel measurements and modeling across diverse sea areas and frequency bands to generalize current findings.
@@
For the minimum resolvable multipath delay interval $\Delta\tau_{2}$, denote with $M$ the number of resolvable MPCs within a time delay bin of width $\Delta\tau_{1}$, then $M\geq 1$. At a certain time instant $t$, the received signal $r_{n}(\tau)$ within the $n$-th delay bin (n=1,,Nn=1,...,N) of width $\Delta\tau_{1}$ can be expressed as [34]